Cancellation law may not hold in an arbitrary ring.


1
Expert's answer
2021-03-17T07:02:38-0400

Let us show that cancellation law may not hold in an arbitrary ring. Let us consider the ring "(\mathbb Z_8,+,\cdot)." Then "\overline{4}\cdot \overline{2}=\overline{4}\cdot \overline{6}=\overline{0}", but "\overline{2}\ne\overline{6}." Therefore, cancellation law does not hold in this ring.